Find the equation of the parabola
a) that is symmetric about the line y=4 and passing through the points (4,4) & (10,0)
b) whose focus and vertex are the vertex and focus, respectively, of the parabola with equation x+4y^2-y=0

Answer by TimothyLamb(4379)

---
This post contains 2 questions
I will answer the first:
---
The general equation of a horizontal parabola with vertex (h,k) is:
x = a(y – k)^2 + h
---
given:
(4,4) and
the parabola is symmetric about the line y=4
---
(4,4) must be the vertex
---
x = a(y – k)^2 + h
x = a(y – 4)^2 + 4
---
given:
(10,0)
---
x = a(y – 4)^2 + 4
10 = a(0 – 4)^2 + 4
10 = a(-4)^2 + 4
10 = 16a + 4
16a = 14
a = 14/16
a = 7/8
---
answer:
x = (7/8)(y – 4)^2 + 4
---
in standard form:
x = (7/8)(y – 4)(y – 4) + 4
x = (7/8)(yy – 8y + 16) + 4
x = (7/8)yy – 7y + 14 + 4
---
x = (7/8)y^2 – 7y + 18
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php